# REFINER: Reasoning Feedback on Intermediate Representations

### v1

@@ 7 Conclusion
In this paper, we propose REFINER, a framework to improve the reasoning abilities of LMs through an iterative feedback loop between two models, a generator and a critic. Our evaluation of this framework on three reasoning tasks showed structured and fine-grained feedback on intermediate reasoning errors results in significant performance gains, surpassing scalar value feedback. Our trained critic model alone, even when noisy, can improve intermediate representations of LMs, showing that REFINER can significantly boost LMs’ performance on reasoning tasks. Our REFINER framework is very general and in principle, might be applied to steer language models in performing different reasoning tasks. More specifically, critic model can be seen as a tool for LLMs to refine their generation quality.
@@ Acknowledgment
Our REFINER framework could not be comprehensively evaluated on all applicable downstream reasoning tasks due to their sheer number. While deliberately distinct, we focused on only three different reasoning tasks in order to study how natural language reasoning feedback can impact downstream tasks. We believe this represents an initial but important step towards exploring automated natural language feedback on intermediate representations. In addition, the critic we presented here is specific for each task, while the ideal critic would be a general one, capable of providing feedback on a wide range of reasoning tasks. Similarly, we considered fine-grained reasoning errors specific to each reasoning task. Recent work has mentioned several other fine-grained reasoning errors (Golovneva et al., 2023), which can’t be fully covered by the reasoning tasks we considered. Generalizing both the critic and fine-grained error types emerge as both the main limitations of this paper and the directions of future work.
@@
Appendix A REFINER Framework
@@
Appendix B Datasets and Models
In Table 6, we report the data statistics. In Table 7, we report the details of used models.
@@ Appendix B Datasets and Models
Appendix C Additional Results
Answer Prediction ($y$) Acc %
GTS 30.8
Graph2Tree 36.5
BERT-Tree 32.4
Roberta-large-GTS 41.0
Roberta-large-Graph2Tree 43.8
Roberta-large-DeductReasoner 45.0
Few-Shot GPT-3 63.05
Few-Shot GPT-3 + COT 63.5
Few-Shot GPT-3 + COT + REFINER 66.4
Table 8: Results on SVAMP dataset
@@ Appendix C Additional Results
Appendix D Qualitative Examples
Figure 8 and 12 depict a qualitative example of REFINER where REFINER could correct incorrect equations through structured feedback, fixing the operators within a multistep solution. Table 12 shows some qualitatively improved examples for MS.
Figure 7: An overview of the three tasks tackled in this paper, with examples of both valid and invalid intermediate reasoning steps, as well as their corresponding fine-grained error types. Notice the Missing Steps error type, in the second task, actually encompasses two error types: reasoning misalignment, derived from not considering the or operation, and lack of implicit knowledge, where implicit knowledge is needed to match the existing rules.
@@ Appendix D Qualitative Examples
Appendix E Feedback Data Generation
Based on the error types as mentioned in Fig 7 and Table 10, we perturb the plausible hypotheses ($z$) in the training data and collect a pool of data $D$ ($x$: input, $z$: plausible hypothesis, $z^{\prime}$: implausible hypothesis). Since our perturbations are based on logic and reasoning errors, we create structured feedback $f$ for every example ($x,z,z^{\prime}$) by stating the error type that occurs in $z^{\prime}$ but not in $z$. For e.g., in sNLR, we omit a few inference steps from the correct hypothesis "#0: viridian is green, #1: rose is green" and create an incorrect (incomplete) hypothesis. For such a scenario, we create feedback "Missing link between fact and rules". In Table 11, we show some examples from the training data for MS.
@@ Appendix E Feedback Data Generation
Appendix F Human Evaluation on Moral Stories
As part of the human evaluation of model generations on MS, we asked Amazon MTurk (AMT) annotators to judge the relevancy of the generated norm and the moral action based on a Likert scale, with 1 = strongly disagree, 2 = disagree, 3 = unsure, 4 = agree, and 5 = strongly agree. Ratings were subsequently aggregated, with scores $\geq$ 4 deemed to be Relevant and with scores, $\leq$ 2 deemed to be Irrelevant while ratings with score 3 (Unsure) left as is. More specifically, we asked three different human judges to evaluate each example. We performed majority voting over answers with the rating Unsure assigned to those examples with no clear majority winner. In Figures 9 and 10, we report a complete breakdown of evaluation results for both norm and moral action. We also report agreement scores computed according to Krippendorff’s $\alpha$ Krippendorff (2018) in Table 4. The low and moderate $\alpha$ values indicate that judging the plausibility of moral norms and actions is a challenging task. In Figures 11-19, we provide excerpts of HIT instructions given to AMT workers during moral norm and action evaluation. Each task was supplemented by an Acceptance and Privacy Policy (Figure 19) that explains participation and data collection terms. All workers were based in US and paid \$0.10 per task which took around 5 minutes to complete on average.

### v2

@@ 7 Conclusion
In this paper, we propose REFINER, a framework to improve the reasoning abilities of LMs through an iterative feedback loop between two models, a generator and a critic. Our evaluation of this framework on three reasoning tasks showed structured and fine-grained feedback on intermediate reasoning errors results in significant performance gains, surpassing scalar value feedback. Our trained critic model alone, even when noisy, can improve intermediate representations of LMs, showing that REFINER can significantly boost LMs’ performance on reasoning tasks. Our REFINER framework is very general and, in principle, might be applied to steer language models in performing different reasoning tasks. More specifically, the critic model can be seen as a tool for LLMs to refine their generation quality.
@@ Acknowledgment
Our REFINER framework could not be comprehensively evaluated on all applicable downstream reasoning tasks due to their sheer number. While deliberately distinct, we focused on only three different reasoning tasks in order to study how natural language reasoning feedback can impact downstream tasks. We believe this represents an initial but important step towards exploring automated natural language feedback on intermediate representations. In addition, the critic we presented here is specific for each task, while the ideal critic would be a general one, capable of providing feedback on a wide range of reasoning tasks. Similarly, we considered fine-grained reasoning errors specific to each reasoning task. Recent work has mentioned several other fine-grained reasoning errors (Golovneva et al., 2023), which can’t be fully covered by the reasoning tasks we considered. Generalizing both the critic and fine-grained error types emerges as both the main limitations of this paper and the directions of future work. Finally, with LLMs being deployed more and more for real-life applications (medical domain, making important decisions), we believe it is crucial to develop expert models and automatic feedback mechanisms to inspect model generations and improve them. LLMs are impressive and work well on several NLP tasks, but they are not expert systems. Our work aims to address this gap by showing that adding interventions/feedback from critics (specialised finetuned critics) can help the LLM model to be more accurate—additionally, making the whole process more transparent.
@@
Appendix A Additional Results
A.1 More details about the quality of the feedback
Please note we also include instances where the critic feedback should say the solution is correct and hence generate ’No’. Our exact match metric is not order-sensitive. We extract the sentences and match them individually to the oracle answers. Since we focused only on the semi-structured critic feedback, automatic evaluation can already capture (measure effectively) the quality of the feedback.
A.2 Details about ReACT and Self-consistency and Self-Correct
The ReACT method consists of the reason model (Reason-Only) LLM (GPT-3.5), which generates a single thought at each step, and the Action model LLM (another GPT-3.5) does the calculation and generates the intermediate outputs (observations). We propose to refine the intermediate steps generated by the above steps and report the results below. Please note ReAct is approx 3-4 times more expensive than GPT-3.5 + CoT. In our experiments, we assumed $3$ reasoning steps for ReACT and a sample size of $5$ for self-consistency to be more cost-effective. Interestingly, we observe that ReACT perform similarly to CoT for the SVAMP dataset. One intuitive reason is that the SVAMP dataset contains questions which require one or two-hop reasoning only. We find that REFINER performs (+2.2) better than Self-correct (Welleck et al., 2023) on the GSM8K dataset, indicating the importance of correcting the intermediate steps can lead to better performance. Please note that we have used GPT-Neo as the generator model and the Unified QA T5-base model as the critic model, consistent with the Self-correct paper by Welleck et al. (2022).
A.3 More results on SVAMP dataset
Model Accuracy
GPT-Neo (1.3B) 8.5
GPT-Neo + Self-Correct 21.2
GPT-Neo + REFINER 23.4 +/- 0.3
Table 8: Comparing REFINER with self-correct on GSM8K dataset
Answer Prediction ($y$) Acc %
GTS 30.8
Graph2Tree 36.5
BERT-Tree 32.4
Roberta-large-GTS 41.0
Roberta-large-Graph2Tree 43.8
Roberta-large-DeductReasoner 45.0
Few-Shot GPT-3 63.05
Few-Shot GPT-3 + COT 63.5
Few-Shot GPT-3 + COT + REFINER 66.4
Table 9: Results on SVAMP dataset
Appendix B REFINER Framework
@@
Appendix C Datasets and Models
In Table 10 and Table 12, we report the data statistics and dataset details. In Table 11, we report the details of the used models. Our research is conducted solely on datasets that are in the English language.
@@ Appendix C Datasets and Models
Dataset/Tools Citation Link License
SVAMP Patel et al. (2021) https://github.com/arkilpatel/SVAMP MIT License
GSM8k Cobbe et al. (2021b) https://github.com/openai/grade-school-math MIT License
sNLR Liang et al. (2022) https://github.com/stanford-crfm/helm Apache License
Moral Norm Emelin et al. (2021) https://github.com/demelin/moral_stories MIT License
HuggingFace Wolf et al. (2020) https://github.com/huggingface/transformers Apache License
Table 12: More details about datasets and Tools
Figure 5: Examples. REFINER on MWP task. There are different scenarios are highlighted in the figure, where (a) the critic model provides correct feedback, generator model utilizes the feedback and fixes the incorrect equation, (b) the critic model provides a correct feedback however, generator model fails to fix the incorrect equation, and (c) the critic model provides an incomplete feedback generator model partially fixes the incorrect equation.
Model Eq. ($z$) Ans. (y{y})
UQA-large 46.7
UQA-large + PPO 48.2
REFINER large 53.8
REFINER large + Oracle (T=3) 68.1
GPT-$3.5$ + CoT 59.3 63.5
GPT-$3.5$ + CoT + REFINERcritic 62.3 66.4
GPT-$3.5^{\star}$ + CoT 64.1 67.1
GPT-$3.5^{\star}$ + CoT + REFINERcritic 67.3 70.6
@@
Figure 6: Feedback Data Generation. The top row illustrates an example from the sNLR task, where the error types are logically invalid, missing links, and missing implicit knowledge steps. The bottom row illustrates an example from moral norm generation, where the error types are contradiction and semantic misalignment. We perturbed used the plausible intermediate steps to implausible.
Appendix D Training Details
Training Details.
For each task, we train a UnifiedQa-T5-base model (UQA-base) (Khashabi et al., 2020) as a critic (§3.1). Further evaluation details are provided in Appendix G. For exploration (§3.2), we use nucleus sampling with $p=0.5$. We select the hyper-parameters by the validation loss: for both the generator and critic model, we use the Adam optimizer with a learning rate of $1e^{-4}$. Each model is trained for $20$ epochs with early stopping based on validation loss. We trained all models on one A100 GPU. We run our models with $3$ random seeds and report the average results. We perform a binomial sign test. We find that p-values are always <0.05 when we compare REFINER with all the baselines (GPT-3.5, Self-refine, Self-reflection), suggesting our results are not random and significant. For the human study, we selected outputs from the best models (baselines and our model) according to automatic metrics. We train models with $T=3$ iterations. We trained the critic model for 8 hours and trained the generator model for 12 hours.
At inference time, we use greedy decoding for the generator and critic model with $T=1$ for the automatic critic and $T=3$ for the oracle critic. We evaluate our methods using the metrics presented in the original papers that proposed the tasks. On the MWP and sNLR tasks, we use the exact match (EM) metric for intermediate steps (equation generation and inference rules) and accuracy (Acc) for the final answers. For MS, we conduct a manual evaluation study to assess the relevance of norms and moral actions. [^7] Since the automatic scores such as BLUE, ROUGE, etc. only account for word level similarity between gold norms or actions and generate norms or actions.
Appendix E Qualitative Examples
Figure 7 and 20 depict a qualitative example of REFINER where REFINER could correct incorrect equations through structured feedback, fixing the operators within a multistep solution. Table 20 shows some qualitatively improved examples for MS.
@@ Appendix E Qualitative Examples
Appendix F Feedback Data Generation
F.1 Rule-based Perturbation
[…]
For MWP and sNLR problems, the underlying reasoning requires symbolic systems with closed-world rules. Hence, we consider a simple rule-based method to automatically generate the pairs of errors and their corresponding structured feedback by considering the error types and position of the errors (see Fig. 6 and Table 1).
In the moral norm generation task, we consider two kinds of fine-grained errors: logical contradiction and semantic misalignment (incoherent, uninformative). Moral norms are people’s subjective judgments about the character and actions mentioned in the context. Each moral norm is a combination of two components (implicit structure): a moral judgment [You shouldn’t] and an action [criticize your family’s religion]. Firstly, to create logical contradictions, we use the concept of deontic logic from Kiehne et al. (2022) and derive new norms contrary to those of Moral Stories. Hence, we replace the correct moral judgments in the plausible hypothesis with inverse judgments. For example, replacing [You shouldn’t] from the plausible hypothesis to [It’s good], as depicted in Fig. 6. To scale such inverse norms (implausible hypothesis), we paraphrase them by substituting the adjectives with synonyms from WordNet. Secondly, to create semantic misalignments, we must collect implausible hypotheses that are either misaligned with the plausible hypothesis or incomplete in nature. To create them, we replace the correct action (verb phrase) from the plausible hypothesis with random verb phrases selected from the context of the plausible hypothesis.
F.2 Synthetic Feedback Generation
We used a few-shot setting where we varied the instruction, the number of demonstrations, and the formatting of the demonstrations. Since data generation with GPT-$3.5$ is expensive, we generated 3030K, 2020K, and 3030K implausible hypotheses for MWP, sNLR and MS tasks, respectively.
Initial PROMPT: Math Word Problem
You are a helpful assistant for math word problems.
We will provide you with a math word problem,
and your task is to generate the intermediate mathematical
equations as a step for solving the problem
and the final correct answer. Here are two examples:
“Question : ” <Problem Statements> Let’s think step by step
<equation> Answer: <answer>
“Question: ” <Problem Statements> Let’s think step by step
<equation> Answer: <answer>
“Question: ” <Problem Statements> Let’s think step by step
Table 16: Prompts used for generating correct answer given a math word problem
REFINEMENT PROMPT: Math Word Problem
You are a helpful assistant for math word problems.
We will provide you with a math word problem
a solution (containing an equation and an answer),
and feedback on the solution.
Your task is to generate a refined intermediate equation
as a step and the final correct answer.
Here are two examples:
“Question : ” <Problem Statements> Let’s think step by step
<equation> Answer: <answer>
Feedback: <feedback> <equation> Answer: <answer>
“Question: ” <Problem Statements> Let’s think step by step
<equation> Answer: <answer>
Feedback: <feedback> <equation> Answer: <answer>
“Question: ” <Problem Statements> Let’s think step by step
<equation> Answer: <answer>
Feedback: <feedback>
Table 17: Prompts used for generating correct answer given a math word problem
PROMPT: Synthetic Incorrect Instance Generation
You are a helpful assistant for
generating counterfactual reasoning steps.
We will provide you with a problem, an error type
and a correct intermediate reasoning step.
Your task is to generate an incorrect reasoning step
based on the error type.
Here are a few examples for each error type:
“Question : ” <Problem Statements> Let’s think step by step
<correct intermediate steps> Error type: <error type>
Counterfactual: <incorrect intermediate steps>
“Question: ” <Problem Statements> Let’s think step by step
<correct intermediate steps> Error type: <error type>
Counterfactual: <incorrect intermediate steps>
“Question: ” <Problem Statements> Let’s think step by step
<correct intermediate steps> Error type: <error type>
Counterfactual:
Table 18: Prompts used for generating synthetic incorrect instances
@@ F.2 Synthetic Feedback Generation
Appendix G Human Evaluation on Moral Stories
As part of the human evaluation of model generations on MS, we asked Amazon MTurk (AMT) annotators to judge the relevancy of the generated norm and the moral action based on a Likert scale, with 1 = strongly disagree, 2 = disagree, 3 = unsure, 4 = agree, and 5 = strongly agree. Ratings were subsequently aggregated, with scores $\geq$ 4 deemed to be Relevant and with scores, $\leq$ 2 deemed to be Irrelevant while ratings with score 3 (Unsure) left as is. More specifically, we asked three different human judges to evaluate each example. We performed majority voting over answers with the rating Unsure assigned to those examples with no clear majority winner. In Figures 8 and 9, we report a complete breakdown of evaluation results for both norm and moral action. We also report agreement scores computed according to Krippendorff’s $\alpha$ Krippendorff (2018) in Table 4. The low and moderate $\alpha$ values indicate that judging the plausibility of moral norms and actions is a challenging task. In Figures 10-18, we provide excerpts of HIT instructions given to AMT workers during moral norm and action evaluation. Each task was supplemented by an Acceptance and Privacy Policy (Figure 18) that explains participation and data collection terms. All workers were based in US and paid \$0.10 per task which took around 5 minutes to complete on average.